Hydrochloric acid is usually purchased in a concentrated form that is 37.0% HCl by mass and has a density of 1.20 g/mL.

How much concentrated solution would you take to prepare 3.05 L of 0.545 M HCl by mixing with water?
Answer #1

Let the volume of solution be 100 ml

D = M / V

1.2 = M / 100

M = 120 g

37% HCl will be (37 x 120) / 100 = 44.4 g

Number of moles of HCl = 44.4 / 36.5 = 1.22

Molarity = no. of moles / volume of solution in litre

M = 1.22 / 0.100

M = 12.2 M

Now using the formula

M1 V1 = M2 V2

12.2 x V1 = 0.545 x 3.05

V1 = 0.13625 L

or 136.25 mL